OpenStudy (ffasinger):

What is the midpoint of the line segment with endpoints (3.2, 2.5) and (1.6, -4.5)? A. (4.8, -1) B. (2.4, -1) C. (2.4, -2) D. (4.8, -2)

OpenStudy (ffasinger):

There is no picture

OpenStudy (ffasinger):

@jabez177

jabez177 (jabez177):

I just got up so I'm not fully in the mood of helping people at the moment. Sorry. :P

OpenStudy (ffasinger):

Ok thanks .

OpenStudy (ffasinger):

@kelsey.benton00

OpenStudy (kelsey.benton00):

yikes, sorry it's been a while since I've dealt with midpoints. @Englishguy is the genius :p also ask maybe @CyberShadow he thinks hes pretty smart

OpenStudy (ffasinger):

Lol thanks

OpenStudy (anonymous):

\[x=\frac{ x _{1}+x _{2} }{ 2 },y=\frac{ y _{1}+y _{2} }{ 2 }\]

OpenStudy (mathmale):

Note that you could sketch this situation yourself, and perhaps even estimate the midpoint from eyeballing y our sketch. Or, to do this problem analytically, use surjithayer's suggested midpoint formulas (which are correct). Pls show your work.

OpenStudy (ffasinger):

@okdutchman7

OpenStudy (okdutchman7):

What is your question?

OpenStudy (mathmale):

So, you're not planning to sketch this situation, or to apply the midpoint formulas given you by surjithayer?

OpenStudy (ffasinger):

At the top @okdutchman7

OpenStudy (ibrafacts):

u should apply those coordination on a graph

OpenStudy (kayders1997):

Use the formula @surjithayer gave you

OpenStudy (kayders1997):

@FFASinger

OpenStudy (ffasinger):

im here

OpenStudy (mathmale):

FFASinger: You are, of course, free to ignore my advice and guidance in regard to solving the problem you've posted, if you wish. But have you stopped to consider how much time you are losing through your apparent wait for someone else to do your work for you? Again, here are the midpoint formulas:\[x=\frac{ x _{1}+x _{2} }{ 2 },~~y=\frac{ y _{1}+y _{2} }{ 2 }\] If you don't know how to use them, then please ask questions.

OpenStudy (anonymous):

B is the correct answer

OpenStudy (kayders1997):

@innocentkyara please don't give dorect answers right away

OpenStudy (anonymous):

X=(3.2+1.6)/2 and Y=(2.5+-4.5)/2 Then (X,Y)=(2.4,-1) Then (x,y)=B